the ratio of girls to boys in a school is 3:7. on a Monday, 1/4 of the girls were absent and 1/6 of the boys were absent. if 16 fewer learners had been absent, then 7/8 of the school would have been present. calculate the number of learners present at school that day


1
Expert's answer
2022-04-28T08:47:49-0400

Let there be "N" learners, "\frac{3}{10}N" girls and "\frac{7}{10}N" boys.

On Monday "\frac{1}{4}\cdot\frac{3}{10}N+\frac{1}{6}\cdot \frac{7}{10}N=\frac{23}{120}N" learners were absent. And "\frac{97}{120}N" learners were present.


We have that "\frac{23}{120}N-16=\frac{1}{8}N"

"\frac{1}{15}N=16"

"N=240"


"\frac{97}{120}N=194"


Answers: 194 learners.